The temperature of 4.5kg of water is raised from 15 degrees to 100 degrees at constant atmospheric pressure. The specific heat capacity of water is 4.18kj/kgK. Determine the heat transfer required


1
Expert's answer
2022-01-12T08:28:51-0500

"Q=cm(t_2-t_1)=4180\cdot 4.5\cdot(100-15)=1.6~MJ."